OpenStudy (anonymous):

if a number decreases from 1,000 to 805 what is the rate of decrease?

OpenStudy (zzr0ck3r):

do we have a time?

OpenStudy (whpalmer4):

there is no rate of decrease, but there is an amount of decrease. To have a rate, you must have a time... If you want to find the percentage change from A to B, calculate: \[\frac{B-A}{A}*100\%\] Positive result means increase, negative result means decrease. For example, going from 10 to 8: \[\frac{8-10}{10}*100\%= -20\%\] Going from 8 to 10:\[\frac{10-8}{8}*100\% = +25\%\]